OpenStudy (anonymous):

Can someone break down the steps for me? x^6 − 511x^3 − 512 = 0

OpenStudy (anonymous):

try this... let u=x^3 can you factor this?: u^2 - 511u -512 = 0 (u - ?? )(u + ????) = 0 can you figure out what replaces the ?? and ????

OpenStudy (anonymous):

-512, 1